OpenStudy (anonymous):

What are the next three terms in the sequence 11, 18, 25, 32, ...? 37, 44, 51 39, 46, 53 41, 48, 55 43, 50, 57

OpenStudy (anonymous):

The equation for the series is a(n) = 7(n)+4

OpenStudy (the_fizicx99):

B

OpenStudy (anonymous):

Yes, B

OpenStudy (anonymous):

A 10-foot board is to be cut into 3 pieces. Two of the pieces will be the same length and one piece will be 2 feet longer than the other two. Which equation could be used to solve for the lengths? x + x + 2x = 10 x + x + x - 2 = 10 x + x + x = 10 + 2 x + x + x + 2 = 10